# Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch. [Experiment, Analysis & Benchmark]

### v1

@@ 4.2. Scalability
Table 8. Single-host scalability with the GAT model using the NVLink host. Times in seconds. (#Layers, hidden size, #heads, fanout): ogbn-arxiv - (3, 1024, 6, 25), Reddit - (2, 1024, 4, 20), ogbn-products - (3, 128, 3, 20), Orkut - (3, 128, 2, 5). ET - Epoch time, TTA - Time to Accuracy.
@@
Table 9. Distributed scalability with the GraphSage model using 4 GPUs per host connected with NVLink . Times in seconds. (#Layers, hidden size, fanout): Orkut - (3, 512, 10), ogbn-products100M - (2, 512, 10). ET - Epoch time, TTA - Time to Accuracy. DistDGL- not cached, PipeGCN-Cached, BNS-GCN-Cached
@@
Next we answer Q2: How do systems scale with a varying number of GPUs and a varying number of hosts? For each dataset, we pick a GNN architecture that fits in a single GPU and use the single-GPU training performance as a basis to calculate the scalability speedups.
4.2.1. Single-host scalability (NVLink)
We now report the single-host scalability of our two approaches by varying the number of GPUs. We consider the GAT model since it is the most computationally demanding among the ones we considered.
@@ 4.2.1. Single-host scalability (NVLink)
Table 8 presents the results from our scalability experiment. The performance trends when training on a single GPU are similar to the ones we observed for the multi-GPU setups we discussed previously. The full-graph training systems have a much lower epoch time than the mini-batch ones but a higher time-to-accuracy.
@@
When scaling to multiple GPUs, one important advantage of mini-batch training systems is that they can keep the amount of work per GPU constant by increasing the mini-batch size. This is not possible in full-graph training systems since the total amount of work per epoch is constant. As the number of GPU increases, the amount of work per GPU decreases but the communication cost increases, making it difficult to achieve scalability.
This effect is clear when comparing systems using synchronous training. With PipeGCN, which uses full-graph training, the cost of synchronous cross-GPU communication during the forward and backward pass of training hides the gains of reducing the computation cost per GPU.
Several optimizations in full-graph training systems address this problem. PipeGCN has a much better scalability when using asynchronous training. This enables PipeGCN to use pipelining, one of its most important features, to overlap communication and computation and to scale to up to 3.4×3.4$\times$3.4 ×. Sampling of boundary nodes in BNS-GCN also enables good scaling, up to 3.5×3.5$\times$3.5 ×, since the cross-GPU communication is avoided. The mini-batch systems scale almost linearly despite using synchronous training: both DGL and Quiver scale by up to 3.9×3.9$\times$3.9 ×. Besides scalability, these systems achieve lower time-to-accuracy that the full-graph training systems in all settings. We only show ET and ET Speed-up for orkut dataset, since it is a SNAP dataset (Yang and Leskovec, 2012) with only graph structure and synthetic input features and target labels.
We also experiment with GraphSAGE model, and the results are shown in the extended version (Bajaj et al., 2024). For GraphSAGE we see that synchronous PipeGCN doesn’t scale at all, asynchronous PipeGCN scales up to $\times3$ and BNS-GCN scales up to 2.7×2.7$\times$2.7 ×. While DGL and Quiver are able to scale up to 4.0$\approx$ 4.0$\times$ 4.0 × and 3.8×3.8$\times$3.8 × respectively.
Takeaway. The mini-batch training systems show better scalability than the full-graph training ones. This is because they can tune the mini-batch size to keep the work per GPU constant to maintain high GPU utilization, unlike the full-graph training systems. The full-graph training systems also experience increasing communication costs as the number of GPUs increases. However, these overheads can be mitigated using asynchronous training and boundary nodes sampling.
4.2.2. Distributed scalability (NVLink)
@@
We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has four GPUs. We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution. We do not consider Quiver because it does not support distributed training.
Table 9 shows the results for distributed scalability with NVLink enabled devices. The synchronous version of PipeGCN cannot scale when increasing the number of hosts. Using asynchronous training and boundary nodes sampling can make full-graph training systems scale much better, as we also observed for the single-host experiments. DistDGL has similar scalability trends despite the fact that it does not use GPU caching. Despite its higher epoch time compared to full-graph training systems, DistDGL achieves a much lower time-to-accuracy.
Takeaway. The general scalability considerations we made for single-host scalability apply also to multi-host scalability. Synchronous full-graph training does not scale to multiple host. By using asynchronous training and boundary nodes sampling, PipeGCN and BNS-GCN achieve similar scalability as DistDGL, which uses mini-batch training. However, DistDGL has a lower time-to-accuracy in all configurations.
4.2.3. Single-host scalability (PCIe hosts)
We run the same experiments on PCIe hosts (as done in Section 4.1.3), and observe similar results, which are shown in the. We scale the number of GPUs from one to four.
@@ 4.2.3. Single-host scalability (PCIe hosts)
All the full-graph training systems have a lower scalability because they require communication over the PCIe bus, which is slower than NVLink. This slowdown compared to the NVLink case is particularly evident for PipeGCN sync, which can only scale up to 2.0×2.0$\times$2.0 ×. PipeGCN async can scale up to at most 3.1×3.1$\times$3.1 × and BNS-GCN up to at most 3.0×3.0$\times$3.0 ×.
Mini-batch training systems have a better scalability because they require little cross-GPU communication. DGL scales up to 3.8×3.8$\times$3.8 × and Quiver up to 3.5×3.5$\times$3.5 ×.
Takeaway. The full-graph systems are more sensitive to the performance of the cross-GPU interconnect because they need to exchange vertex features between partitions at each GNN layer during the forward and backward pass. When GPUs must communicate over a slow PCIe interconnect, the scalability of these systems is lower. The mini-batch training systems use data-parallelism, which does not require exchanging vertex features across GPUs. Therefore, these systems show better scalability in the PCIe host.
@@ 5. Accuracy Evaluation
Even though a system takes longer than another to reach the same target accuracy, it could still be preferable to use it if it can converge to a higher accuracy. Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
@@
Figure 4 summarizes the accuracy values reported in the literature we reviewed for different GNN models and datasets. Each point in the graph represents the test accuracy reported for a specific GNN model-dataset pair in one of the papers we reviewed. The figure illustrates a wide distribution of accuracy values for the same dataset and model combination.
@@
In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the best accuracy we can achieve using full-graph and mini-batch training? Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method? In practice, the cost of training a model also depends on how long it takes to find good hyperparameters. Therefore, we also consider the following question: Q5) How sensitive are the two training methods to changes in the hyperparameters?
@@
Figure 4. Test accuracy reported in literature on GNN training. (Wan et al., 2022b; dgl, 2023; Hu et al., 2020; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Md et al., 2021; Yang et al., 2023; Li et al., 2024; Gasteiger et al., 2022; Mostafa, 2022; Kaler et al., 2022; Yao and Li, 2021; Zou et al., 2019; Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2019; Chen et al., 2018b, a; Yao and Li, 2021; Wu et al., 2024; Veličković et al., 2018; Thorpe et al., 2021; Frasca et al., 2020; Wan et al., 2022a; Kipf and Welling, 2017; Jia et al., 2020)
@@
Methodology To compare maximum accuracies among full-graph and mini-batch training, we use the methodology depicted in Figure 5. Given a dataset and a GNN model, we conduct two separate hyperparameter searches, one using full-graph training and one using mini-batch training. This is because the two training methods can have different sets of best hyperparameters for the same dataset and model. The space of hyperparameters we consider is shown in Table 3. The best hyperparameter settings we obtained using the two training methods are denoted as FG-search and MB-search. We then run full-graph and mini-batch training on both settings, measure the test accuracy, and denote it as FG-train and MB-train respectively. This yields a total of four accuracy combinations per dataset and GNN model. We report all the hyperparameter settings we found in the extended version of this paper (Bajaj et al., 2024).
@@
To factor out any accuracy bias due to systems-level optimizations, we perform synchronous training with PipeGCN and DGL. We validate our results by verifying that given a dataset, model, and training method, our hyperparameter search achieves accuracy values that are equal to or better than the best accuracy value reported in the literature reviewed in Section 2.3.
@@
5.1. Accuracy with the Best Setting
We start by answering question Q3) What is the best accuracy we can achieve using full-graph and mini-batch training? Answering this question shows whether there is one method that consistently reach higher accuracy than another. When the hyperparameters are properly set, whether full-graph training achieves higher accuracy than mini-batch training or vice versa depends on the dataset-GNN model pairs.
The accuracy results for the GraphSAGE model are shown in Table 10. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
The results show that the accuracies are similar across training methods. For the GraphSAGE model, each method can yield higher accuracy than the other, depending on the dataset. The largest gap between the accuracies obtained with the two methods is less than $0.9\%$ in favor of full-graph training and averaging around $0.3\%$. We ran a similar analysis for the GCN and GAT model, and we report the results in the extended version of this paper (Bajaj et al., 2024). With the GAT model, the largest gap is slightly higher, up to $1.1\%$ in favor of full-graph training and the average is around $0.6\%$. Gaps are higher for GCN, which is a simpler model and turns out to be more sensitive to the choice of hyperparameters. The largest gap in this case is $3.6\%$ in favor of mini-batch training and the average is around $1.27\%$.
Takeaway. These results indicate that no training method alone guarantees higher accuracy than the other and the gap between the two methods is typically small (i.e., less than 3.6% accuracy). If achieving the highest accuracy is an absolute priority regardless of the training cost, it is advisable to conduct hyperparameter search twice using both training methods. Otherwise, using either of the two methods is likely to produce similar accuracy in the absence of other systems-level optimizations such as asynchronous training or boundary node sampling.
@@ 5.1. Accuracy with the Best Setting
5.2. Accuracy with Sub-Optimal Setting
We now answer the next question Q4) What accuracy can we achieve when using a training method with the best hyperparameters found for the other method? This shows how the outcome of the comparison between the two training approaches can change when using sub-optimal hyperparameter settings.

### v3

@@ 4.2. Scalability
Table 5. Single-host scalability for GAT (NVLink host). ET - Epoch time (s), TTA - Time to Accuracy (s).
@@
Table 6. Distributed scalability for GraphSage (NVLink hosts). ET - Epoch time (s), TTA - Time to Accuracy (s).
@@
Next, we answer Q2: How do systems scale with a varying number of GPUs and a varying number of hosts? For each dataset, we pick a GNN architecture that fits in a single GPU and use the single GPU training performance to calculate the scalability speedups.
Single-host scalability (NVLink). We now measure single-host scalability by varying the number of GPUs. We report the hyperparameters in the appendix in table 14 For the orkut dataset, we do not report TTA since it has synthetic input features and target labels.
@@
Table 5 presents the results for GAT models. Like we observed previously, the full-graph training systems have a much lower ET than the mini-batch ones but a higher TTA in all configurations.
@@
Mini-batch systems have a lower TTA than full-graph ones also in a single GPU setting, indicating that they have a lower computation cost to start with. The mini-batch systems scale almost linearly, up to 3.9×3.9$\times$3.9 ×. When scaling to multiple GPUs, one important advantage of mini-batch training systems is that they can keep the amount of work per GPU constant by increasing the mini-batch size. This is not possible in full-graph training systems since the total amount of work per epoch is constant. As the number of GPUs increases, the amount of work per GPU decreases but the communication cost increases, hindering scalability. This explains why full-graph training systems that optimize communication scale better than the Full-Graph baseline. PipeGCN has much better scalability because it overlaps communication and computation using asynchronous training, scaling to up to 3.4×3.4$\times$3.4 ×. Sampling of boundary nodes in BNS-GCN also enables good scaling, up to 3.5×3.5$\times$3.5 ×, since the cross-GPU communication is avoided.
Results on GraphSAGE models are shown in the appendix in table 21 The Full-Graph baseline does not scale at all, PipeGCN scales up to $\times3$ and BNS-GCN scales up to 2.7×2.7$\times$2.7 ×. In contrast, DGL and Quiver are able to scale up to 4.0$\approx$ 4.0$\times$ 4.0 × and 3.8×3.8$\times$3.8 × respectively.
@@
Multi-host scalability (NVLink). We now consider scaling the training of larger graphs (orkut and ogbn-papers100m) in a distributed setting of NVLink hosts, where each host has 4 GPUs. For this experiment, we replace DGL with DistDGL, its distributed version. We report the hyperparameters in the appendix in table 15. We consider the GraphSAGE model because it is the only one implemented in the DistDGL distribution. We run Quiver with data parallelism across multiple hosts.
Figure 4. Convergence curve for GraphSage (3 NVLink hosts).
Table 6 and Figure 4 show the results for distributed scalability with NVLink-enabled devices. Also, in this case, we observe that full-graph training systems consistently have lower epoch time but larger time-to-accuracy compared to DistDGL. Going from single-host to multi-host training impacts different systems in a different way. Full-graph training systems must use the network to exchange vertex features across hosts, which is much slower than NVLink. The Full-Graph baseline actually has a slower epoch time than a single-host setup. Optimizing communication using asynchrony (PipeGCN) and boundary node sampling (BNS-GCN) is even more crucial to scaling than in the single-host scalability experiments. DistDGL partitions the dataset across multiple hosts and performs distributed sampling, which introduces an additional communication cost compared to a single-host implementation. When running on multiple hosts, the DistDGL baseline has a much larger epoch time than Quiver because it uses distributed CPU-based sampling and does not use caching.
@@
Scalability with the PCIe host. We run single-host scalability experiments on the PCIe host, from 1 to 4 GPUs. Like for the previous experiments, mini-batch training systems have a lower time-to-accuracy than full-graph training ones despite having a larger epoch time. The results show that all the full-graph training systems have lower scalability since they must not communicate over the slower PCIe bus instead of NVLink. This slowdown compared to the NVLink case is particularly evident for the Full-Graph baseline, which can only scale up to 2.0×2.0$\times$2.0 ×. PipeGCN async can scale up to at most 3.1×3.1$\times$3.1 × and BNS-GCN up to at most 3.0×3.0$\times$3.0 ×. Mini-batch training systems have better scalability because they replicate most graphs in the cache and require less cross-GPU communication. DGL scales up to 3.8×3.8$\times$3.8 × and Quiver up to 3.5×3.5$\times$3.5 ×.
We also run all systems using 3 PCIe hosts for the ogbn-papers100m graph. Compared to the distributed NVLink setup, hosts are connected by a much slower network (1 Gbps instead of 25 Gbps). Like in the previous experiments, DistDGL has a lower time-to-accuracy than the other full-graph training systems, but the difference is smaller than with the distributed NVLink setup because of the cost of distributed sampling and feature loading over a very slow network. DistDGL outperforms the BNS-DGL system by only $12\%$. The distributed sampling implementation of DistDGL is not optimized for very slow network links. The full results are shown in the appendix in table 20.
Memory scalability. In terms of memory usage, full-graph training systems must keep the entire graph and the related training state partitioned in GPU memory. Mini-batch systems only need to store a micro-batch and the related state in the memory of each GPU. In our experiments, full-graph training systems have a larger peak memory utilization in all the configurations we considered, up to 7.4×7.4$\times$7.4 ×, but the gap decreases when using more GPUs (since mini-batch training systems must store more micro-batches in total), deeper GNNs with more layers, or larger sampling fanouts (since each micro-batch is larger).
Takeaway. Mini-batch training systems have lower time-to-accuracy than full-graph training ones for all models, datasets, and hardware configurations we consider. They also have better scalability in most cases. One important scalability advantage for mini-batch training systems is that they can tune the mini-batch size to keep the work per GPU constant. In full-graph training systems, algorithmic optimizations such as asynchrony and sampling are essential to achieve high scalability. The full-graph systems cannot scale well when they use slower cross GPU interconnects because their communication cost is amplified.
@@ 5. Accuracy Evaluation
Even though a system takes longer than another to reach the same target test accuracy, it could still be preferable to use it if it can converge to a higher test accuracy. Indeed, some works on GNN systems assert that full-graph training achieves higher accuracy and that mini-batch training does not guarantee convergence (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
@@
Figure 5 summarizes the test accuracy values reported in the literature we reviewed for different GNN models and datasets. Each point in the graph represents the test accuracy reported for a specific GNN model-dataset pair in one of the papers we reviewed. The figure illustrates a wide distribution of accuracy values for the same dataset and model combination.
@@
In this Section, we shed light onto this apparent contradiction and answer two questions: Q3) What is the test best accuracy we can achieve using full-graph and mini-batch training? Q4) What test accuracy can we achieve when using a training method with the best hyperparameters found for the other method? Different training systems have different optimizations, so Q5: What is the impact of the optimizations on accuracy?
@@
Figure 5. Test accuracy reported in literature on GNN training (Wan et al., 2022b; dgl, 2023; Hu et al., 2020; Song, Xiang and Aws, A and Usa, Runjie and Ma, and Shanghai, Aws and Lab, A and Li, China and Zhang, Muhan and David Wipf, 2021; Md et al., 2021; Yang et al., 2023; Li et al., 2024; Gasteiger et al., 2022; Yao and Li, 2021; Zou et al., 2019; Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2019; Chen et al., 2018b; Yao and Li, 2021; Wu et al., 2024; Veličković et al., 2018; Thorpe et al., 2021; Frasca et al., 2020; Wan et al., 2022a; Kipf and Welling, 2017; Jia et al., 2020) .
@@
Methodology. To compare maximum test accuracies among full-graph and mini-batch training, we use the methodology depicted in Figure 6. Given a dataset and a GNN model, we conduct two separate hyperparameter searches, one using full-graph training and one using mini-batch training. This is because the two training methods can have different sets of best hyperparameters for the same dataset and model, using the method described in Section 3. We consider vanilla baselines for full-graph and mini-batch training: the Full-Graph baseline and DGL with the default Neighborhood Sampling algorithm, respectively. The best hyperparameter settings we obtained using the two training methods are denoted as FG-search and MB-search. We then run full-graph and mini-batch training on both settings, measure the test accuracy, and denote it as FG-train and MB-train respectively. This yields a total of four test accuracy combinations per dataset and GNN model. We report the hyperparameter settings we found in the appendix in table 12.
@@
We validate our results by verifying that given a dataset, model, and training method, our hyperparameter search achieves accuracy values that are equal to or better than the best accuracy value reported in the literature reviewed in Section 2.3.
@@
Table 8. Accuracy (%) impact of different optimizations - GraphSAGE.
Pubmed Ogbn-arxiv Reddit Ogbn-products Ogbn-papers100M
Full-Graph $77.40\pm 0.04$ $71.76\pm 0.16$ $96.94\pm 0.01$ $79.29\pm 0.15$ $62.82\pm 0.01$
PipeGCN $77.40\pm 0.45$ $71.27\pm 0.66$ $96.93\pm 0.11$ $78.58\pm 0.32$ $61.74\pm 0.13$
BNS-GCN $77.20\pm 0.16$ $71.72\pm 0.27$ $96.95\pm 0.08$ $78.40\pm 0.22$ $62.79\pm 0.10$
AdaQP $77.42\pm 0.59$ $63.73\pm 0.07$ $96.85\pm 0.002$ $78.79\pm 0.004$ $62.54\pm 0.00$
NS $78.20\pm 0.32$ $71.67\pm 0.30$ $96.82\pm 0.04$ $78.84\pm 0.12$ $62.77\pm 0.02$
ClusterGCN $79.72\pm 0.22$ $65.72\pm 0.18$ $95.07\pm 0.03$ $79.50\pm 0.17$ $49.79\pm 0.00$
GraphSaint $84.03\pm 0.61$ $73.80\pm 0.13$ $98.40\pm 0.03$ $77.85\pm 0.07$ $50.32\pm 0.27$
Accuracy with the best setting. We start by answering question Q3: What is the best test accuracy we can achieve using vanilla full-graph and mini-batch training? The accuracy results for the GraphSAGE model are shown in Table 7. Given a dataset, we must compare the accuracies obtained for two combinations: (FG-search, FG-train) for full-graph training and (MB-search, MB-train) for mini-batch training. The results are averaged after 5 runs.
The results show that the accuracies are similar across training methods. For the GraphSAGE model, each method can yield higher accuracy than the other, depending on the dataset. Accuracy tends to have lower variance with full-graph training generally than with mini-batch training. The largest gap between the accuracies obtained with the two methods is less than $0.9\%$ in favor of full-graph training and averaging around $0.3\%$. We ran a similar analysis for the GCN and GAT model, and we report the results in the appendix in table 18 and table 19. With the GAT model, the largest gap is slightly higher, up to $1.1\%$ in favor of full-graph training and the average is around $0.6\%$. Gaps are higher for GCN, which is a simpler model and turns out to be more sensitive to the choice of hyperparameters. The largest gap in this case is $3.6\%$ in favor of mini-batch training and the average is around $1.27\%$.
Accuracy with sub-optimal setting. We now answer question Q4: What test accuracy can we achieve when using a vanilla training method with the best hyperparameters found for the other method?